# Scrutinize What We Ignore: Reining In Task Representation Shift Of Context-Based Offline Meta Reinforcement Learning

### v2

@@ 4.3 Practical Implementation
To show the impacts of the task representation shift issue, we set different manners to adjust the context encoder. We denote the given batch size as $bs$ and the accumulation steps of task representation shift as $acc$. Our manners cover 1) $k=2\times bs,acc=1$, 2) $k=3\times bs,acc=1$, 3) $k=1\times bs,acc=2$, and 4) $k=1\times bs,acc=3$. All manners as well as the original manner, namely $k=1\times bs,acc=1$ are performed for 8 different random seeds.
@@ 5 Experiment
Our experiments are conducted to show that the previous optimization framework that ignores the impacts of task representation shift is not sufficient. We hope to illustrate the necessity of reining in the task representation shift, laying the foundation for further research towards better asymptotic performance of COMRL.
@@ 5.2 MAIN RESULTS
Figure 2: Testing returns of different manners to rein in the task representation shift against the original manner concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 2 illustrates the learning curves of our different manners to rein in the task representation shift against the original manner on 20 testing tasks for each benchmark concerning 3 widely adopted objectives for maximizing $I(Z;M)$. We find that these specific objectives hold unanimous performance demonstration on all benchmarks that the previous training framework, namely $k=1\times bs,acc=1$, cannot provide the best asymptotic performance. These objectives focus solely on the optimization target of the task representation, namely the optimal solution of $I(Z;M)$. While they are guaranteed to produce reasonable results, as shown in Section 4.1, ignoring the impacts of task representation shift still hinders the capability of these algorithms to reach their full potential. As shown by Figure 2, even with minor changes to the algorithms, the performance improvements are substantial.
@@ 5.3 Can The Results Show Consistency Across Different Data Qualities?
Figure 3 demonstrates the learning curves of our different manners to rein in the task representation shift against the original manner concerning 3 widely adopted objectives for maximizing $I(Z;M)$ on Ant-Dir. We find that the contrastive-based algorithm fails on the random dataset. Except for this case, the others show consistent results that reining in the task representation shift may lead to better performance improvements.
@@ 6.1 Can The Pretraining Scheme Be Adopted To Achieve Better Performance Improvement?
Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, Assumption 4.5, the monotonic performance improvement condition for pre-training scheme is:
$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (12)
@@
According to Corollary 6.1, pre-training cannot actually achieve monotonic performance improvements as $|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|$ is not equal to 0. As the task representation cannot be changed under the pre-training scheme, compared to training from scratch, this scheme loses some degrees of freedom to improve performance.
Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $I(Z;M)$. Since the cross-entropy based objective is the direct approximation of $I(Z;M)$, holding better performance under the pre-training scheme is also in line with expectation.
This analysis further enhances the importance of considering the impacts stemming from the variation of the task representation in the alternating process. Nevertheless, how to achieve better performance under the pre-training scheme is also an interesting problem.
@@ 7 Conclusion & Limitation
Finding that the performance improvement guarantee of the COMRL training framework remains underexplored, we constructively introduce the return discrepancy scheme into the COMRL setting, proving that the previous optimization framework can be linked with the general RL objective of maximizing the expected return. Furthermore, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation. Based on this finding, we propose a new issue called task representation shift and prove that it is possible to achieve monotonic performance improvements with proper context encoder updates. With minor changes to the previous training framework for reining in the task representation shift, the performance improvements can be substantial. We prospect that deeply exploring the role of task representation shift can make a profound difference in the COMRL setting. Naturally, one limitation is that there may exist more advanced algorithms to control the task representation shift.
@@ 8.2 Missing Proofs
Denote $Z_{mutual}(\cdot|x)$ is the optimal task representation distribution served for maximizing $I(Z;M)$. Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
@@
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of performance difference in COMRL can be formalized as:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (25)
Proof:
@@
$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z_{mutual}}+G^{\theta_{2}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{mutual}}+G^{\theta_{1}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (31)
Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1, Assumption 4.1 and Assumption 4.5 into the above formulation, we can get:
$G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z_{mutual}}+G^{\theta_{2}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{mutual}}+G^{\theta_{1}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (32)
$\geq-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{2}))+\epsilon^{mutual}_{12}$
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1}))$
@@
$\geq-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+\epsilon^{mutual}_{12}$
$-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (34)
@@
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (35)
@@
Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex. Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8 and Assumption 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder with maximizing $I(Z;M)$ from at least extra $k$ samples, where:
@@
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|+|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ (37)
where $Z^{empirical}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $k$ i.i.d samples from $Z_{mutual}(\cdot|x)$ based on $Z_{1}(\cdot|x)$.
@@
$|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|\leq\sqrt{\frac{\alpha}{k}}$ (38)
Now, we move our focus on solving $|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$.
By applying Lemma 8.2, the $L_{1}$ deviation of the empirical distribution $Z^{empirical}_{2}(\cdot|x)$ and true $Z_{mutual}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
$Pr(|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\geq\epsilon)\leq(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (39)
$Pr(|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|<=\epsilon)\geq 1-(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (40)
@@
$|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (41)
@@
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}$ (42)
Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, and recall that $\kappa$ denotes $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
$\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ (43)
@@
Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, Assumption 4.5, the monotonic performance improvement condition for pre-training scheme is:
$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (45)
@@
$=J^{pretrain}(\theta_{2})-J^{mutual}(\theta_{2})+J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})+J^{mutual}(\theta_{1})-J^{pretrain}(\theta_{1})$ (47)
Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1, Assumption 4.1 and Assumption 4.5 into the above formulation, we can get:
$J^{pretrain}(\theta_{2})-J^{mutual}(\theta_{2})+J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})+J^{mutual}(\theta_{1})-J^{pretrain}(\theta_{1})$ (48)
$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{2}))]+\epsilon^{mutual}_{12}$ (49)
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{1}))]$ (50)
$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|+\epsilon^{mutual}_{12}]$ (51)
To get the monotonic performance improvement, if we consider the condition more strictly, we need the term within the expectation of Eq. (51) to be larger than 0. Hence, we can get:
$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{1}{2\kappa L_{z}}\epsilon^{mutual}_{12}=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (52)

### v3

@@ 4.3 Practical Implementation
To show the impacts of the task representation shift issue, we use different settings to adjust the context encoder. We denote the given batch size as $bs$ and the accumulation steps of task representation shift as $acc$. Our settings include 1) $k=2\times bs,acc=1$, 2) $k=3\times bs,acc=1$, 3) $k=1\times bs,acc=2$, and 4) $k=1\times bs,acc=3$. All settings as well as the original setting, namely $k=1\times bs,acc=1$ are performed for 8 different random seeds.
@@ 5 Experiment
Our experiments are conducted to show that the previous optimization framework that ignores the impacts of task representation shift is not sufficient. We hope to illustrate the potential of reining in the task representation shift, laying the foundation for further research towards better performance improvement of COMRL.
@@ 5.2 MAIN RESULTS
Figure 2: Testing returns of different settings to rein in the task representation shift against the original setting concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 2 illustrates the learning curves of our different settings to rein in the task representation shift against the original setting on 20 testing tasks for each benchmark concerning 3 widely adopted objectives for maximizing $I(Z;M)$. We find that these specific objectives hold unanimous performance demonstration on all benchmarks that the previous training framework, namely $k=1\times bs,acc=1$, cannot provide the best asymptotic performance. These objectives focus solely on the optimization target of the task representation, namely the optimal solution of $I(Z;M)$. While they are guaranteed to produce reasonable results, as shown in Section 4.1, ignoring the impacts of task representation shift still hinders the capability of these algorithms to reach their full potential. As shown by Figure 2, even with minor changes to the algorithms, the performance improvements are substantial.
@@ 5.3 Can The Results Show Consistency Across Different Data Qualities?
Figure 3 demonstrates the learning curves of our different settings to rein in the task representation shift against the original setting concerning 3 widely adopted objectives for maximizing $I(Z;M)$ on Ant-Dir. We find that the contrastive-based algorithm fails on the random dataset. Except for this case, the others show consistent results that reining in the task representation shift may lead to better performance improvements.
@@ 6.1 Can The Pretraining Scheme Be Adopted To Achieve Better Performance Improvement?
Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, the monotonic performance improvement condition for pre-training scheme is:
$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (12)
@@
Due to the implementation error and approximation error, $|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|$ can not equal to 0. Hence, according to Corollary 6.1, pre-training cannot actually achieve monotonic performance improvements. As the task representation cannot be changed under the pre-training scheme, compared to training from scratch, this scheme loses some degrees of freedom to improve performance.
Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $Z^{*}(\cdot|x)$. Notice that there exists the decomposition $|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|\leq|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$. Since the cross-entropy based objective is the direct approximation of $I(Z;M)$, holding better performance under the pre-training scheme is also in line with expectation.
This analysis further enhances the importance of considering the impacts of task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ in the alternating process. Nevertheless, how to achieve better performance under the pre-training scheme is also an interesting problem.
@@ 7 Conclusion & Limitation
Finding that the performance improvement guarantee of the COMRL training framework remains under-explored, we constructively provide a theoretical framework to link the training scheme in COMRL with the general RL objective of maximizing the expected return. After scrutinizing this framework, we further find it ignores the variation of the task representation, which weakens the condition necessary for monotonic performance improvements. Based on this finding, we propose a new issue called task representation shift, refine the condition for monotonicity, and prove that monotonic performance improvements can be achieved with appropriate context encoder updates. With minor changes to the previous training framework for reining in the task representation shift, the performance improvements can be substantial. We prospect that deeply exploring the role of task representation shift can make a profound difference in the COMRL setting. Naturally, one limitation is that there may exist more advanced algorithms to control the task representation shift. Additionally, our work focuses on the representation part for monotonic performance improvement, leaving the policy learning part alone. Hence, one direction that merits further research is to design a policy learning algorithm to achieve better policy improvements conditioning on the optimal task representation distribution. We leave these interesting problems to future work.
@@ 8.2 Missing Proofs
Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
@@
Corollary 8.4 (Monotonic performance improvement condition for previous COMRL works).
When meeting Assumption 4.1, the condition for monotonic performance improvement of previous COMRL works is:
$\epsilon^{*}_{12}\triangleq J^{*}(\theta_{2})-J^{*}(\theta_{1})\geq\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z^{*}(\cdot|x)|)$ (25)
Proof:
$J(\theta_{2})-J(\theta_{1})=J(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J(\theta_{1})$ (26)
Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1 and Assumption 4.1 into the above formulation, we can get:
$J(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J(\theta_{1})$ (27)
$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta_{2})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{2}))]+\epsilon^{*}_{12}$ (28)
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta_{1})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{1}))]$ (29)
$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z(\cdot|x)-Z^{*}(\cdot|x)|+\epsilon^{*}_{12}]$ (30)
To get the monotonic performance improvement, we need Eq. (30) to be larger than 0. Hence, we can get:
$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (31)
@@
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1, the lower bound of the performance difference in COMRL can be formalized as:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (32)
Proof: We can introduce the following decomposition
@@
$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z^{*}}+G^{\theta_{2}}_{Z^{*}}-G^{\theta_{1}}_{Z^{*}}+G^{\theta_{1}}_{Z^{*}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (38)
Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1 and Assumption 4.1 into the above formulation, we can get:
$G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z^{*}}+G^{\theta_{2}}_{Z^{*}}-G^{\theta_{1}}_{Z^{*}}+G^{\theta_{1}}_{Z^{*}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (39)
$\geq-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{2}))+\epsilon^{*}_{12}$
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1}))$
@@
$\geq-\kappa L_{z}|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+\epsilon^{*}_{12}$
$-\kappa L_{z}|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (41)
@@
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (42)
@@
Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex. Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent. When meeting Assumption 4.1, 4.7, 4.8 and 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder via maximizing $I(Z;M)$ from at least extra $k$ samples, where:
@@
$|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|\leq|Z_{2}(\cdot|x)-\tilde{Z}_{2}(\cdot|x)|+|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|$ (44)
where $\tilde{Z}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $k$ i.i.d samples from $Z^{*}(\cdot|x)$ based on $Z_{1}(\cdot|x)$.
@@
$|Z_{2}(\cdot|x)-\tilde{Z}_{2}(\cdot|x)|\leq\sqrt{\frac{\alpha}{k}}$ (45)
Now, we move our focus on solving $|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|$.
By applying Lemma 8.2, the $L_{1}$ deviation of the empirical distribution $\tilde{Z}_{2}(\cdot|x)$ and true $Z^{*}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
$Pr(|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|\geq\epsilon)\leq(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (46)
$Pr(|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|<=\epsilon)\geq 1-(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (47)
@@
$|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (48)
@@
$|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}$ (49)
Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$, and recall that $\kappa$ denotes $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
$\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ (50)
@@
Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, the monotonic performance improvement condition for pre-training scheme is:
$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (52)
@@
$=J^{pretrain}(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J^{pretrain}(\theta_{1})$ (54)
Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1 and Assumption 4.1 into the above formulation, we can get:
$J^{pretrain}(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J^{pretrain}(\theta_{1})$ (55)
$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{2}))]+\epsilon^{*}_{12}$ (56)
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{1}))]$ (57)
$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|+\epsilon^{*}_{12}]$ (58)
To get the monotonic performance improvement, we need Eq. (58) to be larger than 0. Hence, we can get:
$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (59)